OpenStudy (anonymous):

Solve x – 2y = 4 for y.

OpenStudy (anonymous):

-2y=4-x y=-(4-x)/2

OpenStudy (anonymous):

@arch1 , eliminate the negative sign

OpenStudy (anonymous):

y=(x-4)/2 by eliminating negativ sign

OpenStudy (anonymous):

y=-x+4?

OpenStudy (anonymous):

noo y=(x-4)/2